# Background risk model in presence of heavy tails under dependence

### v2

@@ Abstract.
In this work we examine two problems on applied probability, which are directly connected with the dependence in presence of heavy tails. The first problem, is related to max-sum equivalence of the randomly weighted sums in bi-variate set up. Introducing a new dependence, called Generalized Tail Asymptotic Independence, we establish the bi-variate max-sum equivalence, under a rather general dependence structure, when the primary random variables follow distributions from the intersection of the dominatedly varying and the long tailed distributions. On base of this max-sum equivalence, we provide a result about the asymptotic behavior of two kinds of ruin probabilities, over a finite time horizon, in a bi-variate renewal risk model, with constant interest rate. The second problem, is related to the asymptotic behavior of the Tail Distortion Risk Measure, in a static portfolio, called Background Risk Model. In opposite to other approaches on this topic, we use a general enough assumption, that is based on multivariate regular variation.
Keywords: Multivariate regular variation; Joint tail behavior; Randomly weighted sums; Tail distortion risk measure; Bi-variate renewal risk model; Interdependence
@@ 1. Introduction
Simultaneously, the dependence modeling among risks, seems to have equally important impact on insurance applications, while keeps the mathematical interest with respect to generalizations either of some independent results, or of some counterexamples in which the independent results can not be generalized. Hence, we observe that the study of dependence models in combination with the presence of heavy tails, present a useful tool in applications and the same time a strong mathematical support.
@@
At first, we understand the interdependence, as structure between two finite sequences of primary random variables, heavy-tailed distributed, $X_{1},\,\ldots,\,X_{n}$ and $Y_{1},\,\ldots,\,Y_{m}$, such that each of them contains dependent components, but simultaneously the two sequences are dependent each other. In this sense, we introduce a new dependence structure, called Generalized Tail Asymptotic Independent, symbollically $GTAI$, in the Definition 2.1 below, and it belongs to the family of second order asymptotic independence. Next, in section 2, we establish the max-sum equivalence of randomly weighted sums, in bi-variate framework, with random weights $\Theta_{1},\,\ldots,\,\Theta_{n}$, $\Delta_{1},\,\ldots,\,\Delta_{m}$, bounded from above, non-negative and non-degenerated to zero, that are arbitrarily dependent each other, and independent of $X_{1},\,\ldots,\,X_{n}$ and $Y_{1},\,\ldots,\,Y_{m}$, under dominatedly varying and long tailed distributions for the primary variables. Namely we establish the asymptotic relation
@@
called randomly weighted sums. The interdependence now, is expressed by the dependence among the components of vectors ${\bf\Theta}=(\Theta_{1},\,\ldots,\,\Theta_{n})$, and ${\bf X}=(X_{1},\,\ldots,\,X_{n})$, and between the vectors ${\bf\Theta}$ and ${\bf X}$. Such dependence structures were studied in the literature under the framework of multivariate regular variation for the distribution of ${\bf X}$. However, in larger classes of heavy tailed distributions for the components of ${\bf X}$, the interdependence effect appears rarely. The reason is, the difficulty to find max-sum equivalence for the randomly weighted sums, namely
@@
as $x\wedge y\rightarrow\infty$. We refer to [12], [32], [47], [53], [57] for papers that introduced relation (1.3), under a variety of dependence structures and several classes of heavy-tailed distributions. In the paper [5], was for first time established relation (1.3), through interdependence for the case $n=2$, in distribution classes larger than regular variation.
@@
In this subsection, after some necessary notations, we introduce the preliminary material for the heavy-tailed distribution classes. Let denote ${\bf x}=(x_{1},\,\ldots,\,x_{n})$, the scalar product $c\,{\bf x}=(c\,x_{1},\,\ldots,\,c\,x_{n})$, $x\wedge y=\min\{x,\,y\}$, $x\vee y=\max\{x,\,y\}$, $x+=x\vee 0$, $x-=(-x\vee 0)$, $\left\lfloor x\right\rfloor$ the integer part of $x$, and $e_{i}$ represents the vector, whose all the components are 00, except the $i$-th, that is $1$. For two positive functions $f$ and $g$ we write $f(x)=O(g(x))$, as $x\to\infty$, if
@@
All the previous asymptotic notations hold for $x\wedge y\rightarrow\infty$, when we have positive, bi-variate functions. For example we write $f(x,\,y)\sim c\,g(x,\,y)$, as $x\wedge y\rightarrow\infty$, with $c\in(0,\,\infty)$, if
@@
$\dfrac{\overline{V}^{n*}(x)}{\overline{V}(x)}\sim n\,,$
as $x\rightarrow\infty$, where $V^{n*}$ is the n-fold convolution of distribution $V$. The classes $\mathcal{S},\,\mathcal{L},\,\mathcal{H}$ were introduced in [17].
@@
for all (or equivalently, for some) $10<t<$ 1. Let us make clear, that $\mathcal{D}\subsetneq\mathcal{H}$ and $\mathcal{D}\not\subseteq\mathcal{S}$, $\mathcal{S}\not\subseteq\mathcal{D}$. However $\mathcal{D}\cap\mathcal{L}=\mathcal{D}\cap\mathcal{S}\subset\mathcal{L}$. Now, we remind some properties of regular variation. A random variable $Z$ with distribution $V$ is regularly varying with index $\alpha>0$ and write $V\in\mathcal{R}_{-\alpha}$ if
$\dfrac{\overline{V}(tx)}{\overline{V}(x)}\rightarrow t^{-\alpha}\,,$
as $x\rightarrow\infty$, uniformly for any $t>0$. The class of distributions with regularly varying tails is contained in $\mathcal{D}\cap\mathcal{L}$ and further the following inclusion is true
@@
see for example in [31]. Let us now consider the limits
@@
$J_{V}^{+}:=\inf\left\{-\dfrac{\ln\overline{V}_{*}(t)}{\ln t}:t>1\right\}\,,\qquad J_{V}^{-}:=\sup\left\{-\dfrac{\ln\overline{V}^{*}(t)}{\ln t}:t>1\right\}\,.$
@@
The class of regularly varying distributions have many closure properties, see for example [31]. One of this properties is the asymptotic behavior of the tail product convolution, which is the popular Breiman’s Theorem. In the [4], and [18] established the following result. If $Z$ and $\Theta$ two independent random variables with distribution of $Z$ from class $\mathcal{R}_{-a}$, for some $\alpha>0$, and $\Theta$ is non-negative, non-degenerated to zero, such that $\mathbf{E}[\Theta^{\alpha+\epsilon}]<\infty$ for some $\epsilon>0$ then $\mathbf{P}(\Theta\,Z>x)\sim\mathbf{E}(\Theta^{\alpha})\mathbf{P}(Z>x)$, as $x\rightarrow\infty$, which further means that the distribution of the product $\Theta\,Z$ belongs to $\mathcal{R}_{-a}$.
Now we can go to the extension of regular variation in random vectors Let ${\bf X}$ be a random vector in the space $[0,\infty]^{n}$. We remind that ${\bf X}$ follows the multivariate regularly varying distribution, symbolically $MRV$, if there exists a function $b\;:\;\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ and a non-degenerated to zero, Radon measure $\mu$, such that for every $\mu$-continuous Borel set $B\subseteq[0,\infty]^{n}\setminus\{{\bf 0}\}$, it holds
$x\,\mathbf{P}\left[\dfrac{{\bf X}}{b(x)}\in B\right]\stackrel{{\scriptstyle v}}{{\rightarrow}}\mu(B)\,,$ (1.4)
where it refers to the vague convergence, as $x\to\infty$ and we write ${\bf X}\in MRV(\alpha,\,b,\,\mu)$. This measure $\mu$ is homogeneous, namely, for any Borel set $B\subseteq[0,\infty]^{n}\setminus\{{\bf 0}\}$ we obtain
@@
$\dfrac{1}{\overline{V}(x)}\mathbf{P}\left[\dfrac{{\bf X}}{x}\in B\right]\stackrel{{\scriptstyle v}}{{\rightarrow}}\nu(B)$
as $x\to\infty$, for a distribution $V\in\mathcal{R}_{-a}$.
@@
The rest of this paper structure is the following. In Section 2 we study one double background risk model, more specifically we study the tail expectation of this model and we give our main result for tail expectation of unequal heavy-tailedness portfolios under dependence.In section 3 we extend the jointly tail behavior of two weighted random sums, given in [34], in the case of random vectors with generalized tail asymptotically independent components. This dependence structure enables the study of the dependence between two random vectors. Additionally we give the ruin probability in bi-dimensional discrete time risk model under the assumptions of the results of this section. In Section 4, some extensions Breiman’s theorem are presented which used to establish asymptotic results of tail distortion risk measure in various background risk models considering the dependence forms between systemic risk factors and losses.
@@ 2. Generalized tail asymptotic independence
with $\Theta_{i},\,\Delta_{j},\;i,\,j\in{\mathbb{N}}$ arbitrarily dependent non-negative random variables, called random weights, and the primary random variables $\{(X_{i},\,Y_{i}),\;i\in{\mathbb{N}}\}$ are such that $X_{i}$ and $Y_{i}$ are SAI (but $X_{i}$ and $Y_{j}$ are independent for any $i\neq j$), with $\mathbf{P}(X_{i}>x)=\overline{F}_{i}(x)\in\mathcal{D}\cap\mathcal{L}$ and $\mathbf{P}(Y_{j}>x)=\overline{G}_{j}(x)\in\mathcal{D}\cap\mathcal{L}$.
The study of the joint behavior of the tail of two randomly weighted sums, provides a realistic framework for the insurance applications, since in most insurance companies are running several portfolios, which are subject to dependence environment, see for relation (2.1) under several heavy-tailed distribution classes and several dependence structures in [34], [11], [42], [51], etc. Indeed, we find mostly two forms of dependence structure. Firstly, the $\{\Theta_{i},\,\Delta_{j}\}$ are arbitrarily dependent, while the $\{(X_{i},\,Y_{i})\}$ are independent random vectors, within each random vector appears some dependence structure. Secondly, the $\{\Theta_{i},\,\Delta_{j}\}$ are arbitrarily dependent, and in each sequence $\{X_{i}\}$ and $\{Y_{i}\}$ appears some dependence structure, while the two sequences $\{X_{i}\}$ and $\{Y_{i}\}$ are independent. In this paper we combine these two approaches, through the following definition.
@@
$\lim_{\min({x_{i},x_{k},y_{j})}\rightarrow\infty}\mathbf{P}(|X_{i}|>x_{i}\;\mid\;X_{k}>x_{k},\;Y_{j}>y_{j})=0\,,$ (2.2)
@@
$\lim_{\min({x_{i},y_{j},y_{k})}\rightarrow\infty}\mathbf{P}(|Y_{j}|>y_{j}\;\mid\;X_{i}>x_{i},\;Y_{k}>y_{k})=0\,,$ (2.3)
@@
This dependence structure allows dependence between $X_{1},\,\ldots,\,X_{n}$, between $Y_{1},\,\ldots,\,Y_{m}$ and dependence between $X_{i}$ and $Y_{j}$ (not only for $i=j$). In this paper we restrict ourselves in the case with $X_{i},\,Y_{i}$ to be SAI dependent for the same $i$, and $X_{i},\,Y_{j}$ independent for any $i\neq j$.
@@
for all $1\leq j\neq k\leq m$. Through (2.4), (2.5) we obtain the definition of tail asymptotic independence of $X_{1},\,\ldots,\,X_{n}$ and $Y_{1},\,\ldots,\,Y_{m}$ respectively, introduced in [23].
As follows from the last remark, if we choose two mutually independent sequences, where each one has tail asymptotic independent terms, then the structure $GTAI$ is satisfied. We present now an example, which contains interdependence among the two sequences and also satisfies the $GTAI$ structure. For sake of simplicity, we restrict ourselves on non-negative random variables, with $n=m=2$.
@@
as $x\wedge y\rightarrow\infty$, for any $1\leq i\neq k\leq n$ with $j=1,\,\ldots,\,m$, with $a(x)>0$ and $a(x)\rightarrow\infty$, $a(x)=o(x)$, as $x\rightarrow\infty$.
@@
with $a(x)\to\infty$, $a(x)=o(x)$, as $x\to\infty$, with $a\in\mathcal{R}_{0}$. Now, let us define the probabilities
@@
$\mathbf{P}(X_{n}(\Theta)>x,\;Y_{m}(\Delta)>y)\leq I_{1}(x,\,y)+I_{2}(x,\,y)+I_{3}(x,\,y)\,,$ (2.6)
so we have to estimate the $I_{1}(x,\,y),\,I_{2}(x,\,y),\,I_{3}(x,\,y)$,
@@
$\leq\sum_{i=1}^{n}\sum_{j=1}^{m}\sum_{1=k\neq i}^{n}\mathbf{P}\left[\Theta_{i}\,X_{i}>\dfrac{x}{m},\;\Delta_{j}\,Y_{j}>\dfrac{y}{m},\;\Theta_{k}\,X_{k}>\dfrac{a(x)}{n}\right]$
@@
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbf{P}(X_{n}(\Theta)>x,\;Y_{m}(\Delta)>y,\;\Theta_{i}X_{i}>x+a(x),\;\Delta_{j}Y_{j}>y+a(y))$
@@
Now applying [34, Lemma 3 (iii)] in (2.1) and we have the following consequence.
@@
Finally, we present one bi-dimensional discrete time risk model. Recently, the discrete time one-dimensional models have attracted attention by researchers, see [37], [52], [54]. On the other hand more researchers study the multivariate risk models, because it is rarely for an insurance company to operate with one line of business, see [13], [14], [16], [27] among others.
We limit us in only two portfolios and discrete time, where $X_{i}$ and $Y_{i}$ depicts the net loss in $i$-th period, in the first and second line of business respectively, $\Theta_{i}$ and $\Delta_{j}$ they continue to be discount factors.
Therefore, the stochastic discounted value of aggregate insurance losses to time $n\in{\mathbb{N}}$ is described by
@@
thence one type of ruin probability is given as
@@
for any $n\in{\mathbb{N}}$, where $x,y$ are initial capitals in each line of business. This ruin probability depicts the probability that both portfolios have been with negative surplus during the $n$ first periods, but not necessarily simultaneously.

### v3

@@ Abstract.
In this paper, we examine two problems on applied probability, which are directly connected with the dependence in presence of heavy tails. The first problem, is related to max-sum equivalence of the randomly weighted sums in bi-variate set up. Introducing a new dependence, called Generalized Tail Asymptotic Independence, we establish the bi-variate max-sum equivalence, under a rather general dependence structure, when the primary random variables follow distributions from the intersection of the dominatedly varying and the long tailed distributions. On base of this max-sum equivalence, we provide a result about the asymptotic behavior of two kinds of ruin probabilities, over a finite time horizon, in a bi-variate renewal risk model, with constant interest rate. The second problem, is related to the asymptotic behavior of the Tail Distortion Risk Measure, in a static portfolio, called Background Risk Model. In opposite to other approaches on this topic, we use a general enough assumption, that is based on multivariate regular variation.
Keywords: Joint tail behavior; Randomly weighted sums; Tail distortion risk measure; Bi-variate renewal risk model; Interdependence; Multivariate regular variation
@@ 1. Introduction
Simultaneously, the dependence modeling among risks, seems to have equally important impact on insurance applications, while keeps the mathematical interest with respect to generalizations either of some independent results, or of some counterexamples in which the independent results can not be generalized. Hence, we observe that the study of dependent models in combination with the presence of heavy tails, present a useful tool in applications and the same time a strong mathematical support.
@@
At first, we understand the interdependence, as structure between two finite sequences of primary random variables, heavy-tailed distributed, $X_{1},\,\ldots,\,X_{n}$ and $Y_{1},\,\ldots,\,Y_{m}$, such that each of them contains dependent components, but simultaneously the two sequences are also dependent each other. In this sense, we introduce a new dependence structure, called Generalized Tail Asymptotic Independent, symbollically $GTAI$, in the Definition 2.1 below, and it belongs to the family of second order asymptotic independence. Next, in section 2, we establish the max-sum equivalence of randomly weighted sums, in bi-variate framework, with random weights $\Theta_{1},\,\ldots,\,\Theta_{n}$, $\Delta_{1},\,\ldots,\,\Delta_{m}$, bounded from above, non-negative and non-degenerated to zero, that are arbitrarily dependent each other, and independent of $X_{1},\,\ldots,\,X_{n}$ and $Y_{1},\,\ldots,\,Y_{m}$, under dominatedly varying and long tailed distributions for the primary variables. Namely we establish the asymptotic relation
@@
called randomly weighted sums. The interdependence now, is expressed by the dependence among the components of vectors ${\bf\Theta}=(\Theta_{1},\,\ldots,\,\Theta_{n})$, and ${\bf X}=(X_{1},\,\ldots,\,X_{n})$, and simultaneously between the vectors ${\bf\Theta}$ and ${\bf X}$. Such dependence structures were studied in the literature under the framework of multivariate regular variation for the distribution of ${\bf X}$. However, in larger classes of heavy tailed distributions for the components of ${\bf X}$, the interdependence effect appears rarely. The reason is, the difficulty to find max-sum equivalence for the randomly weighted sums, namely
@@
as $x\wedge y\rightarrow\infty$. We refer to [12], [32], [47], [53], [57] for papers that studied relation (1.3), under a variety of dependence structures and several classes of heavy-tailed distributions. In the paper [5], was for first time established relation (1.3), through interdependence for the case $n=2$, in distribution classes larger than regular variation.
@@
In this subsection, after some necessary notations, we introduce the preliminary material for the heavy-tailed distribution classes. Let denote ${\bf x}=(x_{1},\,\ldots,\,x_{n})$, the scalar product $c\,{\bf x}=(c\,x_{1},\,\ldots,\,c\,x_{n})$, $x\wedge y=\min\{x,\,y\}$, $x\vee y=\max\{x,\,y\}$, $x^{+}=x\vee 0$, $x-=(-x\vee 0)$, $\left\lfloor x\right\rfloor$ the integer part of $x$, and $e_{i}$ represents the vector, whose all the components are 00, except the $i$-th, that is $1$. With ${\bf 1}_{\{A\}}$ we denote the indicator function on the set $A$. For two positive functions $f$ and $g$ we write $f(x)=O(g(x))$, as $x\to\infty$, if
@@
We write $f(x)\asymp g(x)$, if it holds both $f(x)=O(g(x))$ and $g(x)=O(f(x))$ as $x\to\infty$. All the previous asymptotic notations hold for $x\wedge y\rightarrow\infty$, when we have positive, bi-variate functions. For example we write $f(x,\,y)\sim c\,g(x,\,y)$, as $x\wedge y\rightarrow\infty$, with $c\in(0,\,\infty)$, if
@@
$\lim_{x\to\infty}\dfrac{\overline{V}^{n*}(x)}{\overline{V}(x)}=n\,,$
where $V^{n*}$ is the n-fold convolution of distribution $V$ with itself. The classes $\mathcal{S},\,\mathcal{L},\,\mathcal{H}$ were introduced in [17].
@@
for all (or equivalently, for some) $10<t<$ 1. Let us make clear, that $\mathcal{D}\subsetneq\mathcal{H}$ and $\mathcal{D}\not\subseteq\mathcal{S}$, $\mathcal{S}\not\subseteq\mathcal{D}$. However $\mathcal{D}\cap\mathcal{L}\equiv\mathcal{D}\cap\mathcal{S}\subset\mathcal{S}$. Now, we remind some properties of regular variation. A random variable $Z$ with distribution $V$ is regularly varying with index $\alpha>0$ and write $V\in\mathcal{R}_{-\alpha}$ if it holds
$\lim_{x\to\infty}\dfrac{\overline{V}(tx)}{\overline{V}(x)}=t^{-\alpha}\,,$
for any $t>0$. The class of distributions with regularly varying tails is contained in $\mathcal{D}\cap\mathcal{L}$ and further the following inclusion is true
@@
see for example in [31, Rem. 2.1, p. 21]. Let us now consider the limits
@@
$J_{V}^{+}:=\inf\left\{-\dfrac{\log\overline{V}_{*}(t)}{\log t}:t>1\right\}\,,\qquad J_{V}^{-}:=\sup\left\{-\dfrac{\log\overline{V}^{*}(t)}{\log t}:t>1\right\}\,.$
@@
The class of regularly varying distributions have many closure properties, see for example [31]. One of this properties is the asymptotic behavior of the tail of product convolution, which is the popular Breiman’s Theorem. In the [4], and [18] established the following result.
If $Z$ and $\Theta$ two independent random variables with distribution of $Z$ from class $\mathcal{R}_{-a}$, for some $\alpha>0$, and $\Theta$ is non-negative, non-degenerated to zero, such that $\mathbf{E}[\Theta^{\alpha+\epsilon}]<\infty$ for some $\epsilon>0$ then
$\mathbf{P}(\Theta\,Z>x)\sim\mathbf{E}(\Theta^{\alpha})\mathbf{P}(Z>x)\,,$
as $x\to\infty$, which further means that the distribution of the product $\Theta\,Z$belongs to $\mathcal{R}_{-a}$.
Now we can go to the extension of regular variation in random vectors. Let ${\bf X}$ be a random vector in the space $[0,\infty]^{n}$. We remind that ${\bf X}$ follows the multivariate regularly varying distribution, symbolically $MRV$, if there exists a function $b\;:\;\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ and a non-degenerated to zero, Radon measure $\mu$, such that for every $\mu$-continuous Borel, namely with $\mu(\partial B)=0$ where $\partial B$ represents the border of $B$, set $B\subseteq[0,\infty]^{n}\setminus\{{\bf 0}\}$, it holds
$\lim_{x\to\infty}x\,\mathbf{P}\left[\dfrac{{\bf X}}{b(x)}\in B\right]=\mu(B)\,,$ (1.4)
and we write ${\bf X}\in MRV(\alpha,\,b,\,\mu)$. This measure $\mu$ is homogeneous, namely for any Borel set $B\subseteq[0,\infty]^{n}\setminus\{{\bf 0}\}$ we obtain
@@
$\lim_{x\to\infty}\dfrac{1}{\overline{V}(x)}\mathbf{P}\left[\dfrac{{\bf X}}{x}\in B\right]=\mu(B)$
for a distribution $V\in\mathcal{R}_{-a}$.
@@ 2. Generalized tail asymptotic independence
with $\{\Theta_{i},\,\Delta_{j},\;i,\,j\in{\mathbb{N}}\}$ arbitrarily dependent non-negative random variables, called random weights, and the primary random variables $\{(X_{i},\,Y_{i}),\;i\in{\mathbb{N}}\}$ are such that $X_{i}$ and $Y_{i}$ are SAI (but $X_{i}$ and $Y_{j}$ are independent for any $i\neq j$), with $\mathbf{P}(X_{i}>x)=\overline{F}_{i}(x)\in\mathcal{D}\cap\mathcal{L}$ and $\mathbf{P}(Y_{j}>x)=\overline{G}_{j}(x)\in\mathcal{D}\cap\mathcal{L}$.
The study of the joint behavior of the tail of two randomly weighted sums, provides a realistic framework for the insurance applications, since in most insurance companies are running several portfolios, which are subject to dependence environment, see for relation (2.1) under several heavy-tailed distribution classes and several dependence structures in [11], [34], [42], [51]. Indeed, we find mostly two forms of dependence structure. Firstly, the $\{\Theta_{i},\,\Delta_{j}\}$ are arbitrarily dependent, while the $\{(X_{i},\,Y_{i})\}$ are independent random vectors, within each random vector appears some dependence structure. Secondly, the $\{\Theta_{i},\,\Delta_{j}\}$ are arbitrarily dependent, and in each sequence $\{X_{i}\}$ and $\{Y_{i}\}$ appears some dependence structure, while the two sequences $\{X_{i}\}$ and $\{Y_{i}\}$ are independent. In this paper we combine these two approaches, through the following definition. In next definition, we use the random variables $X_{1},\,\ldots,\,X_{n}$ and $Y_{1},\,\ldots,\,Y_{m}$, that follow distributions with supports, which are not bounded from above.
@@
$\lim_{{x_{i}\wedge x_{k}\wedge y_{j}}\rightarrow\infty}\mathbf{P}(|X_{i}|>x_{i}\;\mid\;X_{k}>x_{k},\;Y_{j}>y_{j})=0\,,$ (2.2)
@@
$\lim_{{x_{i}\wedge y_{j}\wedge y_{k}}\rightarrow\infty}\mathbf{P}(|Y_{j}|>y_{j}\;\mid\;X_{i}>x_{i},\;Y_{k}>y_{k})=0\,,$ (2.3)
@@
This dependence structure allows dependence between $X_{1},\,\ldots,\,X_{n}$, between $Y_{1},\,\ldots,\,Y_{m}$ and dependence between $X_{i}$ and $Y_{j}$ (not only for $i=j$). In this paper we restrict ourselves in the case with $X_{i},\,Y_{i}$ to be SAI dependent for the same $i$, and $X_{i},\,Y_{j}$ independent for any $i\neq j$. We have to note that the $GTAI$ structure belongs to the family of ’second order asymptotic independnce’, that means that the probability of three or more extremal events, is negligible with respect to the probability of two extremal events, namely one in each sequence.
@@
for all $1\leq j\neq k\leq m$. Through (2.4), (2.5) we obtain the definition of tail asymptotic independence of $X_{1},\,\ldots,\,X_{n}$ and $Y_{1},\,\ldots,\,Y_{m}$ respectively, introduced in [23]. We wonder if the results of our paper can be identified using instead of $GTAI$ the $TAI$ over the $X_{1},\,\ldots,\,X_{n},\,Y_{1}\,\ldots,\,Y_{m}$, as we find under similar frame in [9]. The reply is no, because, inspite of the presence of ’interdependence’ in both cases, the $GTAI$ studies second order asymptotic independence events, while the $TAI$ studies only first order asymptotic independence events. Hence, in Theorem 2.1 below, we demonstrate the ’insensitivity’ with respect to dependence in a more extremal event, in comparison with $TAI$ case.
As follows from the last remark, if we choose two mutually independent sequences, where each one has tail asymptotic independent terms, then the structure $GTAI$ is satisfied. We present now two examples, which contains interdependence among the two sequences and also satisfies the $GTAI$ structure. For sake of simplicity, we restrict ourselves on non-negative random variables, with $n=m=2$.
@@
as $x\wedge y\rightarrow\infty$, for any $1\leq i\neq k\leq n$ with $j=1,\,\ldots,\,m$, where $a(x)>0$, is such that $a(x)\rightarrow\infty$, $a(x)=o(x)$.
@@
where $a(x)>0$, is such that $a(x)\rightarrow\infty$, $a(x)=o(x)$ and $a\in\mathcal{R}_{0}$. Now, let us define the probabilities
@@
$\mathbf{P}(X_{n}(\Theta)>x,\;Y_{m}(\Delta)>y)\leq I_{1}(x,\,y)+I_{2}(x,\,y)+I_{3}(x,\,y)\,.$ (2.6)
Therefore, for the upper bound of the probability in the left hand side of (2.6), it remains to estimate the $I_{1}(x,\,y),\,I_{2}(x,\,y),\,I_{3}(x,\,y)$.
@@
$\leq\sum_{i=1}^{n}\sum_{j=1}^{m}\sum_{1=k\neq i}^{n}\mathbf{P}\left[\Theta_{i}\,X_{i}>\dfrac{x}{n},\;\Delta_{j}\,Y_{j}>\dfrac{y}{m},\;\Theta_{k}\,X_{k}>\dfrac{a(x)}{n}\right]$
@@
$\geq\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbf{P}(X_{n}(\Theta)>x,\;Y_{m}(\Delta)>y,\;\Theta_{i}X_{i}>x+a(x),\;\Delta_{j}Y_{j}>y+a(y))$
@@
Let us remind that, $C_{i}>0$ is the constant from $SAI$ condition, on each pair $(X_{i},\,Y_{i})\,,\;i=1,\,\ldots,\,n\wedge m$, see Assumption 2.2. Now applying [34, Lemma 3 (iii)] in Theorem 2.1 and we have the following consequence.
@@
Finally, we present one bi-dimensional discrete time risk model. Recently, the discrete time one-dimensional models have attracted attention by many researchers, see [37], [52], [54]. On the other hand more researchers study the multivariate risk models, because it is rarely for an insurance company to operate with one line of business, see [13], [14], [16], [27] among others.
We limit us in only two portfolios and discrete time, where $X_{i}$ and $Y_{i}$ depicts the net loss in $i$-th period, in the first and second line of business respectively, $\Theta_{i}$ and $\Delta_{i}$ they continue to be discount factors, of the $i$-th period.
Therefore, the stochastic surplus process of insurer at time $n\in{\mathbb{N}}$ is described by
@@
where $x,y$ are initial capitals in each line of business, hence one type of ruin probability is given as
@@
for any $n\in{\mathbb{N}}$. This ruin probability depicts the probability that both portfolios have been with negative surplus during the $n$ first periods, but not necessarily simultaneously.